OpenStudy (anonymous):

Which equation is an identity? 1)11 – (2v + 3) = –2v – 8 2)5w + 8 – w = 6w – 2(w – 4) 3)7m – 2 = 8m + 4 – m 4)8y + 9 = 8y – 3

OpenStudy (ybarrap):

1

OpenStudy (anonymous):

no #2: 5w + 8 – w = 6w – 2(w – 4) 5w-w+8=6w-2w+8 4w+8=4w+8

OpenStudy (ybarrap):

Mandolino is correct